How would I find the slope of 7.25 inches high and 9 inches across? What equation would I start with?
Answer by smmoore(10)

to find the slope you divide 7.25 by 9 which is rise/run.
7.25/9=0.85555556
then put the known values in y=mx+b form.
put 7.25 in for x, .85555556 in for m, and 9 in for x
then you have 7.25=(.8555555*9)+b
7.25=7.25+b
-7.25 -7.25
0=b
lastly put 0 in for b and .85555556 in for m
the final equation is y=.85555556x+0 or y=.85555556x